# A Regret Perspective on Online Selective Generation

### v1

@@ Appendix H Discussion on Selection Efficiency
$L_{T}(\tau^{\text{triv}})-L_{T}(\tau^{\text{non-triv}})\geq\sum_{t=1}^{T}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\right)=T-\sum_{t=1}^{T}a_{t}(\tau^{\text{non-triv}})>0,$ (42)
where $T-\sum_{t=1}^{T}a_{t}(\tau^{\text{non-triv}})>0$ holds due to $\tau^{\text{non-triv}}\notin\mathcal{{H}}^{\text{triv}}$. This implies that, given sufficiently many data points, any $\tau^{\text{non-triv}}$ that satisfies the FDR guarantee incurs at least $T-\sum_{t=1}^{T}a_{t}(\tau^{\text{non-triv}})$ less loss than $\tau^{\text{triv}}$. This is empirically supported by Figure 21, 22, and 25.

### v2

@@ Appendix I Discussion on Selection Efficiency
$L_{T}(\tau^{\text{triv}})-L_{T}(\tau^{\text{non-triv}})\geq\sum_{t=1}^{T}\frac{\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau^{\text{non-triv}})\neq\texttt{IDK}\right)}{1+\lambda}=\frac{T-\sum_{t=1}^{T}a_{t}(\tau^{\text{non-triv}})}{1+\lambda}>0,$ (81)
where $T-\sum_{t=1}^{T}a_{t}(\tau^{\text{non-triv}})>0$ holds due to $\tau^{\text{non-triv}}\notin\mathcal{{H}}^{\text{triv}}$. This implies that, given sufficiently many data points, any $\tau^{\text{non-triv}}$ that satisfies the FDR guarantee incurs at least $T-\sum_{t=1}^{T}a_{t}(\tau^{\text{non-triv}})$ less loss than $\tau^{\text{triv}}$. This is empirically supported by Figure 22, 23, and 28.